OpenStudy (anonymous):

Complete the first 4 steps for graphing the quadratic function. y = -x^2 - 4x - 3

OpenStudy (anonymous):

Firstly, figure out whether it is maximum or minimum. Is it maximum or minimum?

OpenStudy (anonymous):

I don't even know what that means. In my homework, it says nothing about a "maximum" or a "minimum".

OpenStudy (jtvatsim):

this is an algebra question, right? @Biha this is probably different than the Calculus way of graphing. :)

OpenStudy (anonymous):

Yes, this is algebra.

OpenStudy (anonymous):

y = -x^2 - 4x - 3 y = -(x^2 + 4x + 3) y = -(x+3)(x+1) x-intercepts are at x=-3 and x = -1 The vertex is half-way between at x = -2. Is this correct?

OpenStudy (anonymous):

@xo_kansasprincess_xo and @jtvatsim sorry. I was wrong. Actually, I learn the same thing like this in school but with the different concept.

OpenStudy (anonymous):

It's okay!:)

OpenStudy (jtvatsim):

that looks correct so far, btw :)

OpenStudy (anonymous):

Does the parabola open up or down..?

OpenStudy (jtvatsim):

look at the front of the x^2. If there is a negative sign, then the parabola opens downward. If there is a positive (or nothing), then the parabola opens upward. You can remember this by thinking -x^2 means the parabola has a "negative" attitude and is sad (opens down). But a positive x^2 means that the parabola has a "positive" attitude and is happy (opens up).

OpenStudy (anonymous):

That's a good way to put it

OpenStudy (anonymous):

Are those the first four steps? The answer that I posted

OpenStudy (anonymous):

4 steps to graph a quadratic equation (not necessarily in order) 1) Find y-intercept. Make x = 0 -> y = ... 2. Find the 2 x-intercepts. Make y = 0. Solve the quadratic equation -x^2 - 4x - 3 = 0. Since a -b + c = 0, then one real root is (-1) and the other is (-c/a = -3). 3. Find the coordinate of the parabola axis, at x = -b/2a x = 4/(-2) = -2. 4. Find the min or max. Since a is negative, the parabola is downward, there is a max. Replace x = -2 (coordinate of the parabola axis) y max = -4 + 8 - 3 = 1

OpenStudy (jtvatsim):

@xo_kansasprincess_xo Also, FYI "min" or "max" is just another way of saying "vertex" -- all of these words are commonly used to describe parabolas, it depends on your textbook. :)

OpenStudy (jtvatsim):

There are also many different "4 steps" to solve parabolas, use whichever method you were taught. :)

OpenStudy (anonymous):

1.Determine if the parabola opens up or down. 2.Find the line of symmetry. 3.Find the vertex. 4.Use a table of values to plot a couple of points on each side of the line of symmetry. ^That's what the first four are in my lesson

OpenStudy (anonymous):

it kinda similar with what I am going to say . hoho.

OpenStudy (jtvatsim):

OK, you have answer 1 and 2. The line of symmetry is the x = -2 part. You have almost finished step 3, you just need to find the y value that goes with x = -2. For step 4, you can pick any x values that you want, plug them in and see what the y values would be. And plot those points.

OpenStudy (anonymous):

Lol yeah, but I don't know how to do any of that. That's why I posted the question on here.

OpenStudy (jtvatsim):

OK, the part about finding the y values? Is that the problem? :)

OpenStudy (anonymous):

Any of the steps. I didn't actually do the work for those, I found the answers to the first couple steps on the internet because I don't understand how to do this stuff. Plus, I do online school so I don't have any textbooks or anything like that, and it's hard to contact my teacher's.

OpenStudy (jtvatsim):

Oh, OK... that changes things. I thought that the work you posted was yours. :)

OpenStudy (jtvatsim):

hold on just a second, I'm going to find some references. :)

OpenStudy (jtvatsim):

Alright, let's start from the beginning with step 1. Does the parabola open up or down? Do you understand how to determine that based on what we've said so far?

OpenStudy (jtvatsim):

Here's a quick test: Tell us if the parabola opens up or down. A) x^2 + 3x - 2 B) -3x^2 + 4x + 2 C) -x^2 - 3x - 2 D) 7x^2 -5x + 6

OpenStudy (jtvatsim):

@xo_kansasprincess_xo

OpenStudy (jtvatsim):

In case I have to bail, here is a link to a video explaining how to graph quadratics. They use slightly different methods, but it will get the job done I think: http://www.youtube.com/watch?v=mDwN1SqnMRU